OpenStudy (anonymous):

Some birds and squirrels became stuck on an electrical wire after undergoing routine maintenance. There were 15 animals on the wire with a total of 50 animal feet to “un-stick”. How many birds and how many squirrels were stuck on the wire? Set up the problem, explain why you set it up that way, and provide the solution.

OpenStudy (anonymous):

Hmmm. Well squirrels (technically) have four feet and birds have two. We know 15animals had 50 animal feet. So I'd put it as 4(number of feet squirrels have) and get it close to 50. 4x10=40 and 2x5=10. 40+10=50 and 10+5=15

OpenStudy (anonymous):

Ordinarily I'd have done 4x12=48 BUT that would mean we have 13 animals instead of 15.

OpenStudy (anonymous):

To set it up as an equation we put it 4s+2b=50 and b+s=15

OpenStudy (anonymous):

yeah ok thanks my man reall helped alot.

OpenStudy (anonymous):

No prob.